40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 143     Entry time: Fri Oct 21 15:27:15 2016
Author: Gabriele 
Type: Electronics 
Category: Configuration 
Subject: Removed low pass filter in SUM 

I removed the low pass filter in the QPD SUM signal, used for normalization. This reduced a lot the bump at ~20kHz due to laser intensity noise.

I also switched off the 500 Hz high pass filters in the X_NORM and Y_NORM signals.

ELOG V3.1.3-